# Group Orthogonal Low-Rank Adaptation for RGB-T Tracking

### v1

@@ Conclusion
We demonstrate the detailed parameters of different variants in Table 9. Considering the varying demands for model performance and efficiency across different application scenarios, we design two variants, GOLA-B and GOLA-L, to validate the performance of our method under different backbone scales. Specifically, GOLA-B adopts DINOv2-B224 (dinov2) as the backbone to balance performance and computational efficiency, making it suitable for resource-constrained real-time tracking scenarios. GOLA-L selects DINOv2-L224 as the backbone to enhance feature extraction capabilities for addressing more complex multi-modal interference scenarios. The two variants differ only in backbone scale, while maintaining consistent rank decomposition partition strategies and inter-group orthogonal constraint mechanisms to verify the generalizability and stability of our method across different backbone configurations.
@@
GTOT (gtot). The GTOT dataset represents the initial benchmark meticulously crafted for RGB-T tracking. It encompasses 50 precisely paired video sequences, each containing an infrared video and its associated grayscale video. The videos in this dataset were captured across diverse settings, including laboratories, campus thoroughfares, and playgrounds. As such, GTOT provides a comprehensive set of scenarios for evaluating various trackers.
RGBT210 (rgbt210). The RGBT210 dataset contains 210 RGB and infrared video pairs, all precisely annotated. With over 210,000 frames in total, it provides ground-truth annotations for every tracked object. The dataset is instrumental because it identifies 12 distinct attributes like object deformation and fast motion. These attributes enable a nuanced performance evaluation of trackers by considering specific object behaviours and motions.
RGBT234 (rgbt234). The RGBT234 dataset consists of 234 precisely aligned pairs of visible and infrared video sequences spanning around 234,000 frames. Some individual sequences can be as long as approximately 8,000 frames. This dataset significantly broadens the data availability. It also places a strong emphasis on offering detailed annotation. Each entry has up to 12 distinct attributes, enabling more in-depth data analysis.
LasHeR (lasher). The LasHeR dataset is a large-scale RGB-T tracking dataset, encompassing 1,224 pairs of visible and infrared videos, totalling over 734,800 frame pairs. All frame pairs are spatially aligned and manually annotated with bounding boxes, ensuring high-quality and densely annotated data. The dataset encompasses a wide range of object categories, multiple camera perspectives, and diverse environmental factors, including various seasons, weather conditions, and day-night variations, thereby showcasing a high degree of diversity.
@@
More Attribute-based Comparisons. We select the LasHeR, RGBT234, and RGBT210 datasets to conduct a detailed comparison across different attributes. In Table 14, we compare our method with five state-of-the-art trackers, APFNet (apfnet), ViPT (vipt), TBSI (tbsi), CKD (ckd), and STTrack (sttrack). Results indicate that our approach outperforms STTrack on most attributes. To ensure a fairer comparison across model size dimensions, we perform comparisons in Table 15, where we evaluate all versions of our method against SeqTrackv2 (seqtrackv2) on the LasHeR test set. Even when compared with the larger SeqTrackv2, our best variant consistently demonstrates superior performance across most attributes, validating the effectiveness and robustness of our approach. Additionally, in Table 16 and Table 17, we conduct comparisons of our method with ViPT, TBSI, MPLT (mplt), CKD, and STTrack for RGBT234 and RGBT210, respectively. It is noteworthy that our method demonstrates excellent performance across most attributes. This clearly evidences the strong adaptability of the framework we propose, allowing it to effectively tackle the diverse and complex challenges in different scenarios.
@@
Visualization of orthogonality between groups. In Fig. 7, we show the normalized orthogonality of the low-rank matrices of the model across different groups. For LoRA (lora), we utilize the group indices from GOLA to identify the corresponding ranks for visualization. This result confirms that the orthogonality between enhanced groups effectively reduces redundancy, allowing different rank groups to focus on learning complementary feature dimensions. This structured partitioning of the feature space aligns with the fundamental principles of efficient representation learning. When the model is compelled to explore diverse knowledge under orthogonal constraints, it can more fully leverage the expressive potential of low-rank spaces, ultimately translating into performance enhancement.

### v2

@@ Conclusion
We demonstrate the detailed parameters of different variants in Table 9. Considering the varying demands for model performance and efficiency across different application scenarios, we design two variants, GOLA-B and GOLA-L, to validate the performance of our method under different backbone scales. Specifically, GOLA-B adopts DINOv2-B224 (Oquab et al. 2024) as the backbone to balance performance and computational efficiency, making it suitable for resource-constrained real-time tracking scenarios. GOLA-L selects DINOv2-L224 as the backbone to enhance feature extraction capabilities for addressing more complex multi-modal interference scenarios. The two variants differ only in backbone scale, while maintaining consistent rank decomposition partition strategies and inter-group orthogonal constraint mechanisms to verify the generalizability and stability of our method across different backbone configurations.
@@
GTOT (Li et al. 2016). The GTOT dataset represents the initial benchmark meticulously crafted for RGB-T tracking. It encompasses 50 precisely paired video sequences, each containing an infrared video and its associated grayscale video. The videos in this dataset were captured across diverse settings, including laboratories, campus thoroughfares, and playgrounds. As such, GTOT provides a comprehensive set of scenarios for evaluating various trackers.
RGBT210 (Li et al. 2017). The RGBT210 dataset contains 210 RGB and infrared video pairs, all precisely annotated. With over 210,000 frames in total, it provides ground-truth annotations for every tracked object. The dataset is instrumental because it identifies 12 distinct attributes like object deformation and fast motion. These attributes enable a nuanced performance evaluation of trackers by considering specific object behaviours and motions.
RGBT234 (Li et al. 2019). The RGBT234 dataset consists of 234 precisely aligned pairs of visible and infrared video sequences spanning around 234,000 frames. Some individual sequences can be as long as approximately 8,000 frames. This dataset significantly broadens the data availability. It also places a strong emphasis on offering detailed annotation. Each entry has up to 12 distinct attributes, enabling more in-depth data analysis.
LasHeR (Li et al. 2021). The LasHeR dataset is a large-scale RGB-T tracking dataset, encompassing 1,224 pairs of visible and infrared videos, totalling over 734,800 frame pairs. All frame pairs are spatially aligned and manually annotated with bounding boxes, ensuring high-quality and densely annotated data. The dataset encompasses a wide range of object categories, multiple camera perspectives, and diverse environmental factors, including various seasons, weather conditions, and day-night variations, thereby showcasing a high degree of diversity.
@@
More Attribute-based Comparisons. We select the LasHeR, RGBT234, and RGBT210 datasets to conduct a detailed comparison across different attributes. In Table 14, we compare our method with five state-of-the-art trackers, APFNet (Xiao et al. 2022), ViPT (Zhu et al. 2023), TBSI (Hui et al. 2023), CKD (Lu et al. 2024), and STTrack (Hu et al. 2025a). Results indicate that our approach outperforms STTrack on most attributes. To ensure a fairer comparison across model size dimensions, we perform comparisons in Table 15, where we evaluate all versions of our method against SeqTrackv2 (Chen et al. 2023) on the LasHeR test set. Even when compared with the larger SeqTrackv2, our best variant consistently demonstrates superior performance across most attributes, validating the effectiveness and robustness of our approach. Additionally, in Table 16 and Table 17, we conduct comparisons of our method with ViPT, TBSI, MPLT (Luo et al. 2023), CKD, and STTrack for RGBT234 and RGBT210, respectively. It is noteworthy that our method demonstrates excellent performance across most attributes. This clearly evidences the strong adaptability of the framework we propose, allowing it to effectively tackle the diverse and complex challenges in different scenarios.
@@
Visualization of orthogonality between groups. In Fig. 7, we show the normalized orthogonality of the low-rank matrices of the model across different groups. For LoRA (Hu et al. 2022), we utilize the group indices from GOLA to identify the corresponding ranks for visualization. This result confirms that the orthogonality between enhanced groups effectively reduces redundancy, allowing different rank groups to focus on learning complementary feature dimensions. This structured partitioning of the feature space aligns with the fundamental principles of efficient representation learning. When the model is compelled to explore diverse knowledge under orthogonal constraints, it can more fully leverage the expressive potential of low-rank spaces, ultimately translating into performance enhancement.